# A Coreset Selection of Coreset Selection Literature: Introduction and Recent Advances

### v1

@@ IV-A2 GraNd
The benefit of GraNd/EL2N is that it directly relates to the extent to which a sample drives parameter updates. Moreover, they are straightforward ranking metrics once partial training is done. On the other hand, GraNd can be expensive for very large networks or datasets if computed frequently, while EL2N can be imprecise as an approximation in such scenarios. Among all the methods reviewed, we find that GraNd selects the slowest.
@@ IV-B Decision Boundary Based
Whereas scoring-based approaches track loss signals over time, decision boundary-based methods explicitly estimate how close a sample is to the classification boundary [57, 58]. The assumption underlying these methods is that samples near the boundary tend to be more informative.
@@ IV-B1 DeepFool
DeepFool is a geometry-based method that approximates the distance of a sample to the decision boundary by leveraging adversarial perturbations [27, 59]. The key idea is that samples near the decision boundary carry more information for classification than those farther away, making them ideal for coreset selection, as illustrated in Figure 7.
@@
[…]
For coreset selection, we prioritize samples with the smallest perturbation magnitude $\|\mathbf{r}\|_{2}$, as they are closest to the decision boundary and, thus, more informative. The DeepFool-based method ranks all samples by $\|\mathbf{r}\|_{2}$ and selects the top k with the smallest perturbation norms.
In contrast to training-free geometry-based methods that do not account for the actual learned boundary, DeepFool often yields improved performance. However, it requires repeated adversarial steps, which can be expensive and sensitive to the hyperparameters of adversarial routines (e.g., step size).
@@ IV-B2 Contrastive Active Learning
Contrastive Active Learning (CAL) proposes an acquisition strategy that balances uncertainty sampling and coverage sampling by selecting contrastive examples, i.e., data points that are similar in feature space but yield highly divergent model predictions [60, 61, 62]. This approach ensures that selected examples contribute maximally to refining the decision boundary of the model.
@@ IV-C Submodularity Based
A natural arising question is: Why is it not 100% optimal? The reason we cannot achieve exactly $f(\mathcal{S}^{*})$ with a greedy method is that each step picks the locally best option without accounting for global interactions in the dataset. This is known as greedy suboptimality. However, for most practical submodular maximization problems (including coreset selection), the greedy algorithm with a $(1-1/e)$ guarantee is a well-established strong baseline.
@@ IV-C1 FASS
Consider a dataset $\mathcal{T}$ consisting of labeled ($\mathcal{T}_{l}$) and unlabeled ($\mathcal{T}_{u}$) samples: $\mathcal{T}=\mathcal{T}_{l}\cup\mathcal{T}_{u}$. FASS first selects a candidate pool $\mathcal{C}$ of uncertain examples from $\mathcal{T}_{u}$. A typical uncertainty measure is entropy:
@@
$\mathcal{C}=\operatorname*{arg\,max}_{\mathcal{C}\subseteq\mathcal{T}_{u},|\mathcal{C}|=m}\sum_{\mathbf{x}\in\mathcal{C}}H(\mathbf{x})$ (22)
@@
The key benefit of FASS is that it reduces the computational cost by restricting submodular selection to uncertain points, potentially merging the best of both worlds - uncertainty and coverage. However, the computational costs are still considerably higher than those of other coreset selection methods.
@@ IV-C3 SIMILAR
SIMILAR provides a generalized coreset selection framework applicable to rare-class selection, out-of-distribution data filtering, and uncertainty-aware sampling. By leveraging lazy greedy optimization, it maintains a $(1-1/e)$ approximation guarantee while being more scalable than traditional submodular methods. However, its reliance on computing large similarity matrices increases its complexity, making it less efficient for very large datasets unless approximate submodular maximization techniques are used.
@@ IV-D Gradient Matching Based
In contrast to previous approaches, gradient matching focuses on reconstructing the total gradient of the full dataset using only a subset of samples. This ties subset selection directly to the training optimization path.
@@ IV-E Bilevel Optimization Based
Bilevel optimization frameworks explicitly treat coreset selection as an “outer” problem that optimizes model performance on a validation or labeled set, while the “inner” problem performs standard training. Although more computationally complex, these approaches can produce subsets tuned for generalization rather than simply reconstructing training gradients [73, 74, 75].
@@ IV-E1 RETRIEVE
The RETRIEVE method [76] extends gradient-based coreset selection to the Semi-Supervised Learning (SSL) setting [77], where both labeled and unlabeled data must be leveraged effectively. Unlike GradMatch, which formulates coreset selection as a gradient matching problem in a fully supervised setting, RETRIEVE introduces a bilevel optimization framework that selects the most informative unlabeled samples to improve generalization when labeled data are scarce.
Let $\mathcal{T}=\mathcal{T}_{l}\cup\mathcal{T}_{u}$ denote a dataset consisting of a labeled set $\mathcal{T}_{l}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N_{l}}$ and an unlabeled set $\mathcal{T}_{u}=\{\mathbf{x}_{i}\}_{i=1}^{N_{u}}$. The goal of coreset selection in SSL is to construct a compact subset $\mathcal{S}\subset\mathcal{T}_{u}$ such that training on $\mathcal{T}_{l}\cup\mathcal{S}$ results in a model with a similar or superior generalization to training on the full dataset.
@@
Here, $\mathcal{L}_{u}(\mathbf{x};\theta)$ represents the unsupervised loss applied to unlabeled samples, typically a consistency regularization term such as Mean Squared Error (MSE) between different augmentations.
Solving the bilevel optimization directly is computationally intractable. Instead, RETRIEVE employs a one-step gradient-based approximation to estimate the impact of selecting an unlabeled sample $\mathbf{x}\in\mathcal{T}_{u}$ on the labeled set’s loss:
@@
The coreset $\mathcal{S}$ is then selected by ranking samples according to $s(\mathbf{x})$, ensuring that unlabeled samples with high influence on the labeled loss are prioritized.
@@ IV-E2 GLISTER
The key advantage of GLISTER is its focus on “held-out” performance that can improve real-world robustness (e.g., fewer overfitting concerns). However, it requires setting aside a validation set $\mathcal{V}$, reducing the total data for training and, thereby, potentially removing beneficial samples for the final coreset.
Figure 9: Comparison of method runtimes, selection plus training, in minutes (tested on RTXA6000), sorted in ascending order. The histogram highlights the varying computational costs of the different selection strategies, with GraNd being the most time-intensive and random being the fastest.
@@ IV-F Computation Time
TABLE I: Overview of coreset selection methods discussed in this review.
@@
Geometry-based K-Center Greedy Finds maximal coverage under a geometric measure. Suitable for clustering tasks. Computationally expensive for larger datasets.
Geometry-based Herding Greedily selects samples to minimise distance between coreset and dataset centroids. Fast and easy to implement. Poor performance on complex data distributions.
Training-based Score-based Forgetting Counts forgetting events during training and picks samples with highest counts. Easy to implement. Highly non-deterministic.
GraNd Picks samples with maximal contribution to the loss function during training. Easy implementation. Expensive and slow for large datasets.
Decision Boundary-based DeepFool [27, 59] Iteratively finds the smallest perturbation to move samples across a decision boundary. Picks samples with smallest values. Expensive owing to repeated adversarial steps.
CAL Balances uncertainty and diversity by picking contrastive samples. Computationally intensive; Dependent on embedding quality.
Submodularity-based FASS [29] Combines informativeness and representativeness. Filters out uncer Computationally expensive.
@@
Gradient Matching-based CRAIG [70]. Efficiently accelerates gradient descent. Reduces computational cost. Limited to specific optimization algorithms. May not generalize well to all tasks.
@@
V Blind Coreset Selection
Traditional coreset selection methods rely on labeled data to estimate sample importance via loss values, gradient norms, or decision boundary distances. In contrast, blind coreset selection methods aim to approximate these training dynamics without ground-truth annotations.
@@ V-A ELFS
The Estimation of Label-Free Sample Scores (ELFS) framework [79] achieves label-free coreset selection by leveraging pseudo-labeling through deep clustering [80] to approximate informative selection criteria. More formally, given an unlabeled dataset $\mathcal{T}=\{\mathbf{x}_{i}\}_{i=1}^{N}$, ELFS estimates pseudo-labels $\hat{y}_{i}$ by clustering feature embeddings extracted from a pre-trained network $\phi$:
@@
The whole procedure is visualized in Figure 10. By leveraging pseudo-labels, it mimics training-informed, well-known coreset selection methods (e.g., forgetting events, gradient norms) without explicit supervision. However, the effectiveness of ELFS is highly dependent on how well the clustering algorithm assigns pseudo-labels. This problem becomes pronounced in high-dimensional embeddings, where feature separability is challenging, leading to poor clustering performance.
Figure 11: Coreset selection performances on CIFAR-10. Methods grouped by our taxonomy a colored the same, with dashed lines indicating the different methods within the same category. A general observation is that methods that are good for high pruning ratios ($\alpha$) fail to maintain the performance at low pruning ratios and vice versa. We can see a transition between top-performing methods between 5% and 30% of the remaining dataset size. Concrete values can be found in Table II.
@@ V-B ZCore
Zero-shot Core (ZCore) circumvents the need for model-specific supervision or clustering by employing pre-trained Vision-Language Models (VLMs) to encode feature representations [82]. Given an unlabeled dataset $\mathcal{T}$, the method first embeds each sample into a feature space using a pre-trained foundation model $\phi$ (similar to geometry-based methods):
@@
The key advantage of ZCore is that by eliminating the dependence on model-specific training, it can be applied to any unlabeled dataset without additional computational overhead. However, its reliance on pre-trained representations introduces potential domain bias (similar to other training-free methods). Additionally, while ZCore ensures geometric coverage, it does not explicitly account for task-specific importance, which may limit performance in highly specialized applications.
@@ VI Theoretical Remarks
A fundamental question in coreset selection, and through the lens of the broader field of dataset pruning, is whether a smaller, carefully chosen subset $\mathcal{S}$ can match or even surpass the performance of the full dataset $\mathcal{T}$. This challenge ties into neural scaling laws [83, 84, 85, 86], which describe how the generalization error $\epsilon(\mathcal{T})$ typically follows a power-law decay:
@@
where $\beta$ is a dataset-dependent exponent. This suggests that collecting more data improves model performance but with diminishing returns. However, coreset selection and data pruning in general disrupt this traditional scaling, as they introduce selection bias by favoring the most informative samples [87, 88]. Instead of following the standard power law, a well-pruned dataset can achieve:
$\epsilon(\mathcal{S})\propto|\mathcal{S}|^{-\beta{\prime}},$ (44)
where $\beta{\prime}>\beta$.
This means that with the right subset selection, fewer samples can achieve better generalization than an equally large randomly selected subset. The effective information content of a dataset can be defined as:
@@
Sorscher et al. [90] provide a deeper analysis of how pruning affects generalization. Their study reveals that the optimal subset selection strategy depends on the size of the original dataset:
@@
Moreover, their findings indicate that the pruning ratio should increase as the dataset size increases. This is because, in large datasets, much of the information is redundant, and a smaller fraction of harder examples is sufficient for optimal learning.
Figure 12: Coreset selection performances on CIFAR-10 with 10% corrupted labels, which means that 10% of samples have been reassigned to new class labels [91]. Methods grouped by our taxonomy a colored the same, with dashed lines indicating the different methods within the same category. A general observation is that most methods experience a sharp drop in mid-range pruning ratios (10%-50%), especially geometry-based, gradient matching-based methods, and GraNd (gradient norm). Forgetting, CAL, and GraphCut remain more stable in comparison to the other methods. Concrete values can be found in Table III.
For coreset selection, we see indications of these insights by looking at the performances of different methods on CIFAR-10, as shown in Figure 11 and with concrete values in Table II. Methods emphasizing easy samples are outperforming for higher pruning ratios (smaller dataset size) and underperforming for lower ratios and vice versa (see take-over points at around 10% dataset size). However, Zheng et al. [10], and others [92], partially disagree as they found out that always keeping a small fraction of high-density (redundant) is further boosting the performance alongside a hardest-first sampling strategy.
A crucial takeaway from their work is that pruning can break traditional power law scaling. While random subsampling follows the expected decay $\beta$, a well-pruned dataset can achieve an improved scaling exponent $\beta{\prime}$, leading to exponential improvements in generalization error. This means that dataset pruning is not just about compression - it is a powerful tool for enhancing model performance by maximizing information efficiency.
VI-A Outliers
Another interesting aspect of coreset selection, often overlooked, is the role of outliers [93, 94]. The prior handling of outliers can significantly improve the performance of any coreset selection method [95]. Outliers can either provide valuable decision boundary information or introduce noise and perturbations during the training.
We provide in Figure 12 and Table III additional robustness experiments to highlight the importance of outliers. In this experiment, we corrupted 10% of the labels, i.e., reassigning 10% of samples to new class labels [91]. The selection of corrupted labels was repeated during each coreset sampling procedure.
As a result, one can see that most sampling methods get negatively affected (red), especially geometry-based, gradient-matching-based methods, and GraNd (gradient norm). Interestingly, forgetting, CAL, and GraphCut remain more stable in comparison to the other methods. Furthermore, CAL confirms Equation 44, as its $\alpha=10\%$ pruned dataset outperforms the training on the full (corrupted) dataset. Also, forgetting and GraphCut were able to achieve better results for some cases (green).
As this experiment highlights, future work could further explore automated ways prior to coreset selection to detect and filter outliers dynamically. For instance, Zheng et al. [10] leveraged the area under margin [96] for their coverage-based coreset selection, which captures how far a sample is from the decision boundary over multiple training iterations. Other innovations might include classical outlier detection applications.
TABLE II: Coreset selection performances on CIFAR10 with randomly initialized ResNet-18 [44] models. Values derived from DeepCore [15].

### v2

@@ IV-A2 GraNd
The benefit of GraNd/EL2N is that it directly relates to the extent to which a sample drives parameter updates. Moreover, they are straightforward ranking metrics once partial training is done. However, similar to GraNd, extremely high-sensitivity samples (e.g., outliers) can dominate the sampling distribution [5, 27]. This skew risks overfitting and can reduce coverage of the broader data distribution. Among all the methods reviewed, we find that GraNd has the highest runtime.
@@ IV-B Decision Boundary Based
Whereas scoring-based approaches track loss signals over time, decision boundary-based methods explicitly estimate proximity to the classification boundary [158, 26]. These methods assume that samples near the boundary are more informative, i.e., have a higher importance.
@@ IV-B1 DeepFool
DeepFool approximates a sample’s distance to the decision boundary using adversarial perturbations [34, 95]. In this sense, it is a geometry-based selector. The key idea is that samples near the decision boundary carry more information for classification than those farther away, as illustrated in Figure 7.
@@
$\text{ s.t. }\operatorname*{arg\,max}p(\mathbf{x}+\mathbf{r};\theta)\neq\operatorname*{arg\,max}p(\mathbf{x};\theta).$ (13)
For coreset selection, DeepFool prioritizes samples with the smallest perturbation magnitude $\|\mathbf{r}\|_{2}$, since they are closest to the decision boundary. In contrast to training-free geometry-based methods that do not account for the actual learned boundary, DeepFool often yields improved performance. However, it requires repeated adversarial steps, which can be expensive and sensitive to the adversarial hyperparameters (e.g., step size).
@@ IV-B2 Contrastive Active Learning
Contrastive Active Learning (CAL) balances uncertainty sampling and coverage sampling by selecting contrastive examples [91, 22, 29]. These are points that are close in feature space but yield highly divergent model predictions. This approach ensures that selected examples contribute maximally to refining the decision boundary.
@@ IV-C Submodularity Based
A naturally arising question is: Why is it not 100% optimal? The reason we cannot achieve exactly $f(\mathcal{S}^{*})$ with a greedy method is that each step picks the locally best option without accounting for global interactions in the dataset. This is known as greedy suboptimality. However, for most practical submodular maximization problems (including coreset selection), the greedy algorithm with a (11/e)(1-1/e) guarantee is a well-established strong baseline.
@@ IV-C1 FASS
Consider a dataset $\mathcal{T}=\mathcal{T}_{l}\cup\mathcal{T}_{u}$ consisting of labeled ($\mathcal{T}_{l}$) and label-free ($\mathcal{T}_{u}$) samples. FASS first selects a candidate pool $\mathcal{C}$ of uncertain examples from $\mathcal{T}_{u}$. A typical uncertainty measure is entropy:
@@
$\mathcal{C}=\operatorname*{arg\,max}_{\mathcal{C}\subseteq\mathcal{T}_{u},|\mathcal{C}|=m}\sum_{\mathbf{x}\in\mathcal{C}}H(\mathbf{x}).$ (22)
@@
The key benefit of FASS is that it reduces the computational cost by restricting submodular selection to uncertain points, potentially merging the best of both worlds (uncertainty and coverage). Yet, the computational costs are still considerably higher than those of other coreset selection methods.
@@ IV-C3 SIMILAR
SIMILAR provides a generalized coreset selection framework applicable to rare-class selection, out-of-distribution data filtering, and uncertainty-aware sampling. By leveraging lazy greedy optimization, it maintains a (11/e)(1-1/e) approximation guarantee while being more scalable than traditional submodular methods. However, its reliance on computing large similarity matrices increases its complexity, making it less efficient for very large datasets unless approximate submodular maximization techniques are used.
@@ IV-D Gradient Matching Based
In contrast to previous approaches, gradient matching focuses on reconstructing the total gradient of the full dataset using only a subset of samples. This ties coreset selection directly to the training optimization path.
@@ IV-D2 GradMatch
In summary, gradient-matching methods are training-surrogate selectors: they choose $\mathcal{S}$ so that one (or a few) gradient steps on $\mathcal{S}$ behaves like a gradient step on $\mathcal{T}$, i.e., they approximate the training update direction. In practice, this means they mainly require access to per-sample gradients under the current model and do not need a held-out validation signal. As a result, they often excel at speeding up optimization (fewer examples per step) but are not explicitly optimized for downstream generalization. Any generalization gains are indirect and depend on how well training-gradient reconstruction correlates with validation performance.
@@ IV-E Bilevel Optimization Based
Bilevel optimization frameworks treat coreset selection as model selection for generalization: the subset $\mathcal{S}$ is chosen to directly improve performance on a separate outer objective (typically validation loss or labeled-set loss), while the model parameters are obtained by standard training on $\mathcal{S}$ in the inner problem. This makes bilevel methods practically different from gradient matching in two ways. First, the selection signal is not “does $\mathcal{S}$ reproduce the training gradient?”, but “does training on $\mathcal{S}$ reduce a held-out/generalization loss?”. Second, they typically require either (i) an explicit validation split, or (ii) a labeled subset whose loss acts as a proxy for generalization (as in semi-supervised setups), and they use hypergradient-style approximations (e.g., one-step or truncated inner training) to score candidates. Consequently, bilevel methods tend to be more robust to redundancy and can target distribution shift or noise via the outer objective, but they are more sensitive to how the validation/proxy set is constructed and incur higher overhead due to repeated proxy training [160, 176, 152].
@@ IV-E1 RETRIEVE
The RETRIEVE method [70] extends gradient-based coreset selection to the Semi-Supervised Learning (SSL) setting [121], where both labeled and label-free data must be leveraged effectively. Unlike GradMatch, which formulates coreset selection as a gradient matching problem in a fully supervised setting, RETRIEVE introduces a bilevel optimization framework that selects the most informative label-free samples to improve generalization when labeled data are scarce.
Let $\mathcal{T}=\mathcal{T}_{l}\cup\mathcal{T}_{u}$ denote a dataset consisting of a labeled set $\mathcal{T}_{l}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N_{l}}$ and an label-free set $\mathcal{T}_{u}=\{\mathbf{x}_{i}\}_{i=1}^{N_{u}}$. The goal of coreset selection in SSL is to construct a compact subset $\mathcal{S}\subset\mathcal{T}_{u}$ such that training on $\mathcal{T}_{l}\cup\mathcal{S}$ results in a model with a similar or superior generalization to training on the full dataset.
@@
Here, $\mathcal{L}_{u}(\mathbf{x};\theta)$ represents the unsupervised loss applied to label-free samples, typically a consistency regularization term such as Mean Squared Error (MSE) between different augmentations.
Solving the bilevel optimization directly is computationally intractable. Instead, RETRIEVE employs a one-step gradient-based approximation to estimate the impact of selecting an label-free sample $\mathbf{x}\in\mathcal{T}_{u}$ on the labeled set’s loss:
@@
The coreset $\mathcal{S}$ is then selected by ranking samples according to $s(\mathbf{x})$, ensuring that label-free samples with high influence on the labeled loss are prioritized.
@@ IV-E2 GLISTER
The key advantage of GLISTER is its focus on “held-out” performance that can improve real-world robustness (e.g., fewer overfitting concerns). However, it requires setting aside a validation set $\mathcal{V}$, reducing the total data for training and, thereby, potentially removing beneficial samples.
Figure 9: Comparison of method runtimes, selection plus training, in minutes (tested on RTX A6000), sorted in ascending order. The histogram shows the varying computational costs of the different selection strategies.
@@ IV-F Computation Time
TABLE III: Overview of coreset selection methods discussed in this review.
@@
Geometry-based K-Center Greedy [122] Finds maximal coverage under a geometric measure. Suitable for clustering tasks. Computationally expensive for larger datasets.
Geometry-based Herding [150] Greedily selects samples to minimize distance between coreset and dataset centroids. Fast and easy to implement. Poor performance on complex data distributions.
Training-based Score-based Forgetting [139] Counts forgetting events during training and picks samples with highest counts. Easy to implement. Highly non-deterministic.
GraNd [110] Picks samples with maximal contribution to the loss function during training. Easy implementation. Expensive and slow for large datasets.
Decision Boundary-based DeepFool [34] Iteratively finds the smallest perturbation to move samples across a decision boundary. Picks samples with smallest values. Expensive owing to repeated adversarial steps.
CAL [91] Balances uncertainty and diversity by picking contrastive samples. Computationally intensive; Dependent on embedding quality.
Submodularity-based FASS [149] Combines informativeness and representativeness. Filters out uncertain samples. Computationally expensive.
@@
Gradient Matching-based CRAIG [94] Efficiently accelerates gradient descent. Reduces computational cost. Limited to specific optimization algorithms. May not generalize well to all tasks.
@@
TABLE IV: Hyperparameter sensitivity of representative coreset selectors. We list common sensitive knobs and typical failure modes when mis-set.
Method Sensitive knobs (examples) Typical failure mode
GraNd/EL2N Checkpoint/epoch for scoring; scoring frequency; gradient batch size; seed averaging Early: noisy/hard bias. Late: score collapse or outlier dominance. High variance across seeds.
CAL Encoder/embedding quality; neighborhood size $k$; similarity threshold $\epsilon$; KL/uncertainty weighting; temperature Near-duplicate selection (too strict) or drift to pure uncertainty sampling (too weak similarity constraint); unstable under embedding/domain shift.
GradMatch OMP iterations/budget; regularization $\lambda$; gradient refresh frequency; layer choice; mini- vs full-batch gradients Myopic greedy picks and missed coverage; noisy/oscillatory selections under gradient noise; high runtime if gradients are recomputed too often.
Submodular Similarity metric; kernel bandwidth/temperature; candidate pool size; lazy-greedy settings; coverage vs diversity weighting Kernel miscalibration yields duplicates (too peaky) or overspread subsets (too flat); memory/time blow-up from dense similarity matrices.
DeepFool Max iterations; overshoot/step size; norm choice ($\ell_{2}$ vs $\ell_{\infty}$); stopping criterion; initialization Non-convergence (underestimates boundary proximity) or overshoot (breaks ranking); high compute and sensitivity to calibration.
GLISTER Validation split size/quality; proxy inner steps; proxy LR/regularization; selection batch size Overfits to small/biased validation; unstable gains when proxy training is too short; brittle subsets across seeds/architectures.
V Label-free Coreset Selection
Traditional coreset selection methods rely on labeled data to estimate sample importance via loss values, gradient norms, or decision boundary distances. In contrast, label-free coreset selection methods aim to approximate these training dynamics without ground-truth annotations.
@@ V-A ELFS
The Estimation of Label-Free Sample Scores (ELFS) framework [174] achieves label-free coreset selection by leveraging pseudo-labeling through deep clustering [118] to approximate informative selection criteria. More formally, given an label-free dataset $\mathcal{T}=\{\mathbf{x}_{i}\}_{i=1}^{N}$, ELFS estimates pseudo-labels $\hat{y}_{i}$ by clustering feature embeddings extracted from a pre-trained network $\phi$:
@@
The whole procedure is visualized in Figure 10, namely first applying deep clustering and then classical coreset selection methods afterward. By leveraging pseudo-labels, it mimics training-informed, well-known coreset selection methods (e.g., forgetting events, gradient norms) without explicit supervision. However, the effectiveness of ELFS is highly dependent on how well the clustering algorithm assigns pseudo-labels. This problem becomes pronounced in high-dimensional embeddings, where feature separability is challenging, leading to poor clustering performance.
Figure 11: Coreset selection performances on CIFAR-10. Methods grouped by our taxonomy a colored the same, with dashed lines indicating the different methods within the same category. A general observation is that methods that are good for high pruning ratios ($\alpha$) fail to maintain the performance at low pruning ratios and vice versa. We can see a transition between top-performing methods between 5% and 30% of the remaining dataset size. Concrete values can be found in Table V. Values are transcribed from DeepCore [43].
@@ V-B ZCore
Zero-shot Core (ZCore) circumvents the need for model-specific supervision or clustering by employing pre-trained Vision-Language Models (VLMs) to encode feature representations [42]. Given an label-free dataset $\mathcal{T}$, the method first embeds each sample into a feature space using a pre-trained foundation model $\phi$ (similar to geometry-based methods):
@@
The key advantage of ZCore is that by eliminating the dependence on model-specific training, it can be applied to any label-free dataset without additional computational overhead. However, its reliance on pre-trained representations introduces potential domain bias (similar to other training-free methods).
Additionally, while ZCore ensures geometric coverage, it does not explicitly account for task-specific importance, which may limit performance in highly specialized applications.
V-C Other Notable Mentions
Some very recent works attempt to combine diversity with difficulty without labels. For example, D2Pruning (Diversity & Difficulty Pruning) [89] constructs a graph of data points (nodes) with edges linking similar samples, then spreads “difficulty” scores across this graph to select a subset of examples that are both diverse and hard. They report improved performance over prior methods at moderate pruning rates by capturing complementary aspects of data utility. Another example is SimCore [71], which targets an open-set self-supervised learning scenario by selecting from a large unlabeled pool those samples most similar to a given target dataset’s domain. While SimCore is slightly tangential (aimed at data augmentation for representation learning), it reinforces the trend of using semantic similarity in feature space.
@@ VI Theoretical Remarks
A fundamental question in coreset selection is whether a carefully chosen subset $\mathcal{S}$ can match or even surpass the performance of the full dataset $\mathcal{T}$. This question also sits within the broader area of dataset pruning. This challenge connects to neural scaling laws [63, 50, 25, 40]. These laws suggest that the generalization error $\epsilon(\mathcal{T})$ typically follows a power-law decay:
@@
where $\beta$ is a dataset-dependent exponent. This suggests that collecting more data improves model performance but with diminishing returns. However, coreset selection (and pruning more broadly) can disrupt this traditional scaling. The reason is selection bias: pruning favors the most informative samples [55, 141]. As a result, a well-pruned dataset may deviate from the standard power law:
$\epsilon(\mathcal{S})\propto|\mathcal{S}|^{-\beta{\prime}},\quad\text{ where }\beta{\prime}>\beta.$ (44)
Thus, the right selection can generalize better than a random subset of the same size. The effective information content of a dataset can be defined as:
@@
Sorscher et al. [131] provide a deeper analysis of how pruning affects generalization. They argue that the optimal subset selection strategy depends on the size of the original dataset:
@@
They further argue that the pruning ratio should increase with dataset size. In large datasets, much of the information is redundant, so a smaller fraction of harder examples can suffice. We observe similar trends on CIFAR-10 (see Figure 11 and Table V). These results are consistent with the above insights. Easy-sample strategies perform better at high pruning ratios.
They often underperform at lower ratios, and the reverse holds for hard-sample strategies. We observe take-over points around a 10% dataset size. However, Zheng et al. [173] and others [151] partially disagree. They report that retaining a small fraction of redundant samples can further improve performance when combined with hardest-first sampling.
TABLE V: Coreset selection performances on CIFAR10 with randomly initialized ResNet-18 [47] models. Values are transcribed from DeepCore [43].